Please answer all sub-parts to this one question

Part A - Audible wavelengths. The range of audible
Answer #1

\lambda=\frac{v}{f}

speed of sound = 343.59 metres per second

\lambda_{min}=\frac{ 343.59 m/s}{2*10^4Hz}=17.179*10^{-3}m

\lambda_{max}=\frac{ 343.59 m/s}{20Hz}=17.179m

-------------------------

f=\frac{v}{\lambda}

f_{min}\frac{3*10^8m/s}{700*10^{-9}m}=4.285*10^{14}hz

f_{max}\frac{3*10^8m/s}{400*10^{-9}m}=7.5*10^{14}hz

--------------------------------------

\lambda_s=\frac{ 343.59m/s}{28.5*10^{3}Hz}=1.2125 cm

--------------------------------------

\lambda_s=\frac{ 1480m/s}{28.5*10^{3}Hz}=5.21 cm